Hello I am helping my son on a homework assignment. Unfortunately am stuck on this question. Can someone help me out. Any assistance would be greatly appreciated. Here is the question:

Mary counted her dimes when she put then in groups of 4 she have 2 dimes leftovers.
When she put them in groups of 5 she have 1 leftovers, what is the smallest numbers of dimes she could have if she have more that 10 dimes?

Answer by Earlsdon(6294)

You solve this problem by making two tables:
The first table is for the groups of four with two leftover.
3 X 4 + 2 = 14
4 X 4 + 2 = 18
5 X 4 + 2 = 22
6 X 4 + 2 = 26
The second table is for groups of five with one leftover.
2 X 5 + 1 = 11
3 X 5 + 1 = 16
4 X 5 + 1 = 21
5 X 5 + 1 = 26
The last entries in the two tables show that Mary must have had at least 26 dimes.
She could make 6 groups of 4 dimes with 2 leftover or she could make 5 groups of 5 dimes with 1 leftover.